University of Colorado at Boulder University of Colorado CU Home Search A to Z Index Map
Laboratory for Atmospheric and Space Physics


Polar spacecraft anomaly:

On August 2 (Day 214/20:26), an anomaly occurred in the primary  Command and
Attitude Processor (CAP).  The problem manifested itself in the inability
of the spacecraft to respond to normal ground commands. To regain control,
a soft boot of the CAP was executed by a direct command to the processor,
bypassing the normal command path. Resetting the CAP required the reloading
into memory of all spacecraft and instrument macros, as well as the stored
command table.

Following restoration of the CAP and regaining valid telemetry, the power
system mode controller was found to have switched from the primary to the
secondary side; the system was restored manually to the primary side.
Evaluation of spacecraft and instrument telemetry showed that  main
power had not been interrupted to spacecraft subsystems or instruments,
despite the automatic switch to the secondary power system mode controller.
However, normal operation of the despun platform was interrupted, the
battery V/T levels were changed, and the spacecraft clock was reset to
1995. In addition, the HYDRA instrument showed both high and low voltage
systems to be in the off state;  the TIDE high voltage was also in the off
state.  In both these cases, the main power remained on.

Immediately after restoring command capability, the  clock was reset
manually to GMT, and the battery V/T levels reset to nominal.  The VIS
instrument scientist  elected to place the instrument in a safe state.
In addition, the UVI filter wheel was placed in the "home" position to
close the optical path;  the instrument remained in operation.  CAMMICE,
CEPPAD, MFE, PIXIE, SEPS, EFI, and TIMAS reported the instruments to be
operating normally following the anomaly. On August 5, the  despun platform
was reconfigured for normal operation and calibrated.  HYDRA was restored
to operation on August 4, TIDE on August 5, and VIS on August 6. The UVI
filter wheel was set for normal  operation on August 5, following
reconfiguration of the despun platform.

A subsequent report from EFI indicated that the instrument was in an unusual
state.  A power off reset was executed on August  4 to return EFI to a known
state before being configured for normal operation. The CEPPAD IPS sensor
was later reported to have been reset to the default Table 0 mode; the
normal operating mode uses Table 5.

The spacecraft and all instruments are now in full operation, with no
lasting effects evident.  NOAA's Space Environments Center reported a
large solar flare occurring close to the time of the anomaly. We presume
this to be the cause of the upset of the spacecraft subsystems.

John Wainwright